# Faster Video Diffusion with Trainable Sparse Attention

### v1

@@ 1 Introduction
To support VSA’s hierarchical sparse attention, we prototype GPU kernel where the coarse stage kernel fuses softmax, Top-$\mathcal{K}$ selection, and block indexing into a single pass, and the fine stage leverages a block-sparse attention kernel based on FA dao2022flashattentionfastmemoryefficientexact . This leads to our VSA implementation that retains $85\%$ of FA3’s MFU shah2024flashattention . We further retrofit VSA into SoTA open source DiT, Wan2.1 1.3B wan2025wan , originally trained with full attention. This integration speedups attention time by 6x and reduces end-to-end inference latency from 31s to 18s (1.7x) on H100. As a result, VSA enables the first video DiT where attention accounts for only 20% of runtime during both training and inference without quality degradation.
To our best knowledge, VSA is the first trainable sparse attention approach that, based on extensive experiments totaling around 90k H200 hours, shows better scaling than full attention on DiTs. Finally, we hope that our explicit ablation of the key parameters (e.g., tile size, critial token prediction, locality prior, sparsity, etc.) will enable more targeted exploration of sparse attention in scaling video DiTs.
@@ 3.1 Ablation Studies
Finegrained Attention Map Leads to Better Performance But a Slower Kernel. As analyzed in Section 2.1, the tile size $B$ in VSA is a critical hyperparameter that balances computational efficiency against model performance. It directly affects two key aspects: (1) how accurately critical-tokens can be identified through the coarse stage’s cube size, and (2) the granularity of attention in the fine stage. Hardware constraints partially dictate this parameter—NVIDIA Hopper GPUs optimize for matrix multiplications with dimensions divisible by 16, and smaller tiles generally reduce arithmetic intensity. Our benchmarks in Table 1(d) show that decreasing tile size from $256\times 256$ to $64\times 16$ significantly reduces Model FLOPS Utilization (MFU).
Training with various tile sizes (Table 1(c)) reveals smaller tiles consistently reduce model loss through finer attention granularity. This improvement stems from the finer granularity allowing the coarse stage to more accurately predict critical-token cubes and the fine stage to focus attention on smaller, more relevant regions. Experiment 16 specifically tests mismatched granularity between stages, confirming both coarse and fine stage granularity significantly impact performance. Here, the coarse stage used smaller pooling cubes $(C_{t},C_{h},C_{w})=(4,4,4)$ (effectively $B=64$) while the fine stage operated on larger tiles corresponding to $(C_{t},C_{h},C_{w})=(4,8,8)$ (effectively $B=256$). To reconcile the finer granularity of the coarse stage’s predicted attention map with the coarser block-sparse attention in the fine stage, we applied an additional $(1,2,2)$ average pooling before selecting top-$\mathcal{K}$ entries.
@@
Mean Pooling Is Sufficient. We also examined different pooling methods for the coarse stage. Table 1(e) shows average pooling outperforms both max pooling and convolutional approaches, with the latter causing training instability.
@@ 3.2 Scaling Studies
To validate VSA, we pretrained a 410M video DiT with with latent shape $(16,32,32)$ (16,384 tokens), larger than our 120M ablation models. Figure 2(a) shows VSA achieves nearly identical loss to full attention despite 87.5% sparsity ($\mathcal{K}=32$ out of 256 cubes), while reducing attention FLOPS by 8×8$\times$8 × and end-to-end training FLOPS by 2.53×2.53$\times$2.53 ×. Further scaling experiments from 60M to 1.4B parameters (Figure 2(b)) confirm that VSA consistently produces a better Pareto frontier than full attention. The parallel fitted curves indicate that VSA maintains its 2.53×2.53$\times$2.53 × FLOPS reduction across scales. Each model was trained with compute budgets up to $4\times 10^{21}$ FLOPS on 128 H200 GPUs with sequence length of 16K. To our knowledge, VSA is the first trainable sparse attention for video DiTs demonstrating superior performance compared to full attention under rigorous scaling evaluation. While we leave comprehensive scaling studies at longer sequence lengths to future work, our fine-tuned Wan 2.1 already shows 1.7$\times$ inference speedup at 23K sequence length (§3.3).
@@ 3.3 Sparse Adaptation
To evaluate VSA’s effectiveness, we further finetune Wan-1.3B with VSA on synthetic data generated by Wan-14B with video latent $16\times 28\times 52$ We set $\mathcal{K}$ to 32, corresponding to a 91.2% attention sparsity. As shown in Table 2, VSA achieves even higher VBench huang2024vbench score compared to the original Wan-1.3B. We hypothesize training with synthetic data from a larger model may contributes to this boost. To ensure a fair comparison, we also finetune Wan-1.3B using the same synthetic data. The results show that all models perform closely on VBench, indicating that VSA can retain generation quality despite significant attention sparsity. We additionally compared VSA to SVG xi2025sparsevideo , a training-free attention sparsification method, under extreme sparsity. Figure 3 shows that VSA is preferred even though it has a higher sparsity, demonstrating the effectiveness of training with sparse attention. The DiT inference time of Wan-1.3B drops from 31s (full attention with torch compile) to 18s with VSA.
@@ 3.4 Kernel Performance
As Figure 4(b) shows, VSA’s fine block sparse kernel approaches theoretical limit with nearly 7× speedup over FlashAttention-3 at long sequence lengths (85% MFU over FA3). Even after accounting for the coarse stage computations,  VSA still maintains over 6× speedup. In contrast, FlexAttention dong2024flex with an identical block-sparse mask (64×64 block size) achieves only a 2× speedup. Applying VSA’s speedup to Wan-1.3B and Hunyuan brings 2-3$\times$×inference speedup, as shown in Figure 4(a).
@@ 5 Related Work
Sparse Attention in LLMs. There had been a proliferation of fixed-pattern sparse attention mechanisms in large language models child2019generating ; zaheer2020big ; beltagy2020longformer ; ding2023longnet ; guo2021longt5 . In practice, however, most LLM training (over 90% of total FLOPs) happens on short sequences ($\leq$32K tokens) under the “train-short, adapt-long” paradigm xiong2023effectivelong ; liu2024deepseek ; grattafiori2024llama ; yang2024qwen2 , so sparse attention saw little uptake beyond sliding window attention variants like in Mistral jiang2023mistral7b . With LLMs targeting contexts beyond 1M tokens, sparse attention has seen renewed interest. Recent work primarily targets inference-time speedups  xiao2023efficient ; zhang2023h2o ; jiang2024minference ; xu2025xattention , while the latest methods explore trainable, dynamic sparsity patterns (MoBA lu2025moba , NSA yuan2025native ) to enable efficient end-to-end training on extreme-length sequences. We draw inspirations from them, however, VSA differs from MoBA by directly contributing the coarse-grained attention output to the final representation and using smaller blocks compatible with efficient block-sparse kernels. Compared to NSA, the nature of video and bidirectional attention avoids grouped query constraints of attention pattern. We discuss similarities and difference in depth in §D.
Sparse Attention in Video DiTs. Recent work has explored applying sparse attention post-hoc to DiTs pretrained with full attention zhang2025fastvideo ; ding2025efficient ; zhang2025sparge ; xi2025sparsevideo ; hassani2025generalized at inference. However, we argue that the case for trainable sparse attention in video DiTs is both distinct from and more urgent than in LLMs. First, video DiT demands far longer sequences, e.g., a 100K-token context yields only 5s video, making scaling inherently more costly than. Second, unlike LLMs, where long-context adapaption is a small fraction of total training, state-of-the-art video DiTs kong2024hunyuanvideo ; wan2025wan ; polyak2025moviegencastmedia ; tan2025dsv dedicate most of their compute budget to full-resolution, long-sequence training. As a result, these models remain bottlenecked by quadratic attention at both training and inference. This calls for trainable sparse attention mechanisms, like VSA, as a core design of video DiTs, not a post-hoc fix. DSV tan2025dsv also explore adding sparsity to attention during DiT training, but their multi-stage and profiler-based design may complicate the training pipeline, which we further discuss in §D.
@@ Appendix B Experimental Details
We document the detailed experiments setups for results presented in Section 3.
@@ Appendix C Coarse Stage Runtime
For shorter sequence lengths, the coarse stage overhead is more pronounced. Our profiling experiments using nsys (Table 4) reveals that Top-$\mathcal{K}$ selection dominates this runtime. Although we fused the kernels for attention scaling, softmax, and Top-$\mathcal{K}$ operations to reduce memory traffic and improve data locality, it only provided modest improvements. Since the coarse stage overhead becomes negligible at longer sequence lengths – our primary target – we did not pursue further optimizations in this work. However, coarse stage acceleration remains an important direction for future research.
@@ Appendix E Broader Impact
VSA aims to make high-quality video generation more accessible by significantly reducing the training and inference cost of video diffusion models. Our work may help democratize video creation tools for a broader range of users and use cases, including education, animation, and independent media production. However, as with many generative technologies, VSA also presents potential risks. In particular, the ability to generate realistic videos at scale may increase the risk of malicious applications, such as generating deepfakes or misleading media content. We emphasize the importance of developing robust detection tools, usage guidelines, and ethical standards in parallel with technical advances to ensure responsible deployment of such models.

### v2

@@ 1 Introduction
To support VSA’s hierarchical sparse attention, we prototype GPU kernel where the coarse stage kernel fuses softmax, Top-$\mathcal{K}$ selection, and block indexing into a single pass, and the fine stage leverages a block-sparse attention kernel based on FA dao2022flashattentionfastmemoryefficientexact . This leads to our VSA implementation that retains $85\%$ of FA3’s MFU shah2024flashattention . We further retrofit VSA into SoTA open source DiT, Wan2.1 1.3B wan2025wan , originally trained with full attention. This integration speeds up attention time by 6x and reduces end-to-end inference latency from 31s to 18s (1.7x) on H100. As a result, VSA enables the first video DiT where attention accounts for only 20% of runtime during both training and inference without quality degradation.
To the best of our knowledge, VSA is the first trainable sparse attention approach that, based on extensive experiments totaling around 90k H200 hours, shows better scaling than full attention on DiTs. Finally, we hope that our explicit ablation of the key parameters (e.g., tile size, critical token prediction, locality prior, sparsity, etc.) will enable more targeted exploration of sparse attention in scaling video DiTs.
@@ 3.1 Ablation Studies
Finegrained Attention Map Leads to Better Performance But a Slower Kernel. As analyzed in Section 2.1, the tile size $B$ in VSA is a critical hyperparameter that balances computational efficiency against model performance. It directly affects two key aspects: (1) how accurately critical tokens can be identified through the coarse stage’s cube size, and (2) the granularity of attention in the fine stage. Hardware constraints partially dictate this parameter—NVIDIA Hopper GPUs optimize for matrix multiplications with dimensions divisible by 16, and smaller tiles generally reduce arithmetic intensity. Our benchmarks in Table 1(d) show that decreasing tile size from $256\times 256$ to $64\times 16$ significantly reduces Model FLOPS Utilization (MFU).
Training with various tile sizes (Table 1(c)) reveals smaller tiles consistently reduce model loss through finer attention granularity. This improvement stems from the finer granularity allowing the coarse stage to more accurately predict critical-token cubes and the fine stage to focus attention on smaller, more relevant regions. Experiment 16 specifically tests mismatched granularity between stages, confirming that both coarse and fine stage granularity significantly impact performance. Here, the coarse stage used smaller pooling cubes $(C_{t},C_{h},C_{w})=(4,4,4)$ (effectively $B=64$) while the fine stage operated on larger tiles corresponding to $(C_{t},C_{h},C_{w})=(4,8,8)$ (effectively $B=256$). To reconcile the finer granularity of the coarse stage’s predicted attention map with the coarser block-sparse attention in the fine stage, we applied an additional $(1,2,2)$ average pooling before selecting top-$\mathcal{K}$ entries.
@@
Mean Pooling Is Sufficient. We also examined different pooling methods for the coarse stage. Table 1(e) shows that average pooling outperforms both max pooling and convolutional approaches, with the latter causing training instability.
@@ 3.2 Scaling Studies
To validate VSA, we pretrained a 410M video DiT with latent shape $(16,32,32)$ (16,384 tokens), larger than our 120M ablation models. Figure 2(a) shows VSA achieves nearly identical loss to full attention despite 87.5% sparsity ($\mathcal{K}=32$ out of 256 cubes), while reducing attention FLOPS by 8×8$\times$8 × and end-to-end training FLOPS by 2.53×2.53$\times$2.53 ×. Further scaling experiments from 60M to 1.4B parameters (Figure 2(b)) confirm that VSA consistently produces a better Pareto frontier than full attention. The parallel fitted curves indicate that VSA maintains its 2.53×2.53$\times$2.53 × FLOPS reduction across scales. Each model was trained with compute budgets up to $4\times 10^{21}$ FLOPS on 128 H200 GPUs with sequence length of 16K. To our knowledge, VSA is the first trainable sparse attention for video DiTs demonstrating superior performance compared to full attention under rigorous scaling evaluation. While we leave comprehensive scaling studies at longer sequence lengths to future work, our fine-tuned Wan 2.1 already shows 1.7$\times$ inference speedup at 23K sequence length (§3.3).
@@ 3.3 Sparse Adaptation
To evaluate VSA’s effectiveness, we further finetune Wan-1.3B with VSA on synthetic data generated by Wan-14B with video latent $16\times 28\times 52$ We set $\mathcal{K}$ to 32, corresponding to a 91.2% attention sparsity. As shown in Table 2, VSA achieves even higher VBench huang2024vbench score compared to the original Wan-1.3B. We hypothesize training with synthetic data from a larger model may contribute to this boost. To ensure a fair comparison, we also finetune Wan-1.3B using the same synthetic data. The results show that all models perform closely on VBench, indicating that VSA can retain generation quality despite significant attention sparsity. We additionally compared VSA to SVG xi2025sparsevideo , a training-free attention sparsification method, under extreme sparsity. Figure 3 shows that VSA is preferred even though it has a higher sparsity, demonstrating the effectiveness of training with sparse attention. The DiT inference time of Wan-1.3B drops from 31s (full attention with torch compile) to 18s with VSA.
@@ 3.4 Kernel Performance
As Figure 4(b) shows, VSA’s fine block sparse kernel approaches the theoretical limit with nearly 7× speedup over FlashAttention-3 at long sequence lengths (85% MFU over FA3). Even after accounting for the coarse stage computations,  VSA still maintains over 6× speedup. In contrast, FlexAttention dong2024flex with an identical block-sparse mask (64×64 block size) achieves only a 2× speedup. Applying VSA’s speedup to Wan-1.3B and Hunyuan brings 2-3$\times$×inference speedup, as shown in Figure 4(a).
@@ 5 Related Work
Sparse Attention in LLMs. There has been a proliferation of fixed-pattern sparse attention mechanisms in large language models child2019generating ; zaheer2020big ; beltagy2020longformer ; ding2023longnet ; guo2021longt5 . In practice, however, most LLM training (over 90% of total FLOPs) happens on short sequences ($\leq$32K tokens) under the “train-short, adapt-long” paradigm xiong2023effectivelong ; liu2024deepseek ; grattafiori2024llama ; yang2024qwen2 , so sparse attention saw little uptake beyond sliding window attention variants like in Mistral jiang2023mistral7b . With LLMs targeting contexts beyond 1M tokens, sparse attention has seen renewed interest. Recent work primarily targets inference-time speedups  xiao2023efficient ; zhang2023h2o ; jiang2024minference ; xu2025xattention , while the latest methods explore trainable, dynamic sparsity patterns (MoBA lu2025moba , NSA yuan2025native ) to enable efficient end-to-end training on extreme-length sequences. We draw inspiration from them; However, VSA differs from MoBA by directly contributing the coarse-grained attention output to the final representation and using smaller blocks compatible with efficient block-sparse kernels. Compared to NSA, the nature of video and bidirectional attention avoids grouped query constraints of the attention pattern. We discuss similarities and differences in depth in §D.
Sparse Attention in Video DiTs. Recent work has explored applying sparse attention post-hoc to DiTs pretrained with full attention zhang2025fastvideo ; ding2025efficient ; zhang2025sparge ; xi2025sparsevideo ; hassani2025generalized at inference. However, we argue that the case for trainable sparse attention in video DiTs is both distinct from and more urgent than in LLMs. First, video DiT demands far longer sequences, e.g., a 100K-token context yields only 5s video, making scaling inherently more costly than language models. Second, unlike LLMs, where long-context adaptation is a small fraction of total training, state-of-the-art video DiTs kong2024hunyuanvideo ; wan2025wan ; polyak2025moviegencastmedia ; tan2025dsv dedicate most of their compute budget to full-resolution, long-sequence training. As a result, these models remain bottlenecked by quadratic attention at both training and inference. This calls for trainable sparse attention mechanisms, like VSA, as a core design of video DiTs, not a post-hoc fix. DSV tan2025dsv also explores adding sparsity to attention during DiT training, but their multi-stage and profiler-based design may complicate the training pipeline, which we further discuss in §D.
@@ Appendix B Experimental Details
We document the detailed experiments setups for the results presented in Section 3.
@@ Appendix C Coarse Stage Runtime
For shorter sequence lengths, the coarse stage overhead is more pronounced. Our profiling experiments using nsys (Table 4) reveal that Top-$\mathcal{K}$ selection dominates this runtime. Although we fused the kernels for attention scaling, softmax, and Top-$\mathcal{K}$ operations to reduce memory traffic and improve data locality, it only provided modest improvements. Since the coarse stage overhead becomes negligible at longer sequence lengths – our primary target – we did not pursue further optimizations in this work. However, coarse stage acceleration remains an important direction for future research.
@@ Appendix E Broader Impact
VSA aims to make high-quality video generation more accessible by significantly reducing the training and inference cost of video diffusion models. Our work may help democratize video creation tools for a broader range of users and use cases, including education, animation, and independent media production. However, as with many generative technologies, VSA also presents potential risks. In particular, the ability to generate realistic videos at scale may increase the risk of malicious applications, such as generating deepfakes or misleading media content. We emphasize the importance of developing robust detection tools, usage guidelines, and ethical standards in parallel with technical advances to ensure the responsible deployment of such models.